# Ubiquity of the spin-orbit induced magnon nonreciprocity in ultrathin ferromagnets

### v1

@@ III.2 The case of 2Co/W(110)
Finally it is worth to comment on the presence of the antisymmetric magnetic exchange interaction, known as Dzyaloshinskii–Moriya interaction (DMI) in the investigated systems. Similar to many other ultrathin ferromagnetic films grown on heavy-element substrates, DMI is also active in the present systems. This fact should, in principle, lead to an asymmetric magnon dispersion relation and a nonreciprocity of the magnon energies [41, 42, 49]. However, in the case of Co double layer and Co/Ni bilayer grown on Ir(001) due to the presence of a chirality-inverted DMI the nonreciprocity in the energy is rather small (3 meV or less) [20, 22]. In the case of Co double layer on W(110) the DMI-induced magnon nonreciprocity in the energy is slightly larger [29]. However, the spin-orbit induced magnon nonreciprocity as discussed here is a different phenomenon and appears in the excitation process.
IV Conclusion
@@
The results clearly demonstrate that both the incident beam energy and the choice of the substrate provide ways to tune the magnon nonreciprocity. This may inspire novel ideas for excitation of nonreciprocal ultrafast THz magnons for the application in the field of magnonics. We anticipate that the effect discussed here is not restricted to electron scattering experiments. It should also be observed in transport-based excitation schemes [50, 51, 52, 53]. Moreover, it is not limited to ultrathin films of $3d$ ferromagnets grown on metallic substrates. A similar effect is also expected when the underlying substrate is not metallic , e.g., dielectric oxides [54]. As long as the spin-orbit scattering is large one should be able to observe a magnon nonreciprocity. Analogously, the same effect is also expected for two-dimensional (2D) van der Waals magnets as well as heterostructures based on such 2D materials. One idea is to fabricate a bilayer made of a 2D ferromagnet and another 2D material which possesses a large SOC. Likewise, one may think of a 2D ferromagnet put on top of a topological insulator, e.g., Bi2Se3(0001) [55]. Since the interfacial electronic hybridization is not decisive for this phenomenon, one should, in principle, observe a notable magnon nonreciprocity.
@@ Acknowledgements.
The financial support of the Deutschen Forschungsgemeinschaft (DFG) through the DFG Grants ZA 902/5-1, ZA 902/7-1 and ZA 902/8-1 is acknowledged. We would like to thank the Physikalisches Institut for hosting the group and for providing the infrastructure.

### v2

@@ III.2 The case of 2Co/W(110)
Finally it is worth to comment on the presence of the antisymmetric DMI in the investigated systems. Similar to many other ultrathin ferromagnetic films grown on heavy-element substrates, DMI is also active in the present systems [33]. This fact should, in principle, lead to an asymmetric magnon dispersion relation and a nonreciprocity of the magnon energy (eigenfrequency) [15, 23, 32]. However, as discussed earlier in the case of Co double layer and Co/Ni bilayer grown on Ir(001), due to the presence of a chirality-inverted DMI, the nonreciprocity in the magnon energy is rather small [37, 33]. In the case of Co double layer on W(110) the DMI-induced magnon nonreciprocity in the energy is slightly larger [45]. However, the spin-orbit induced magnon nonreciprocity discussed here is a different phenomenon and appears in the magnon amplitude. This type of nonreciprocity becomes more significant at smaller magnon wavevectors (see Figs. 4 and 6).
IV Concluding remarks
@@
The fact that the magnon nonreciprocity for the case of 2Co/W(110) presented in Fig. 9 is smaller (compared to that of the 2Co/Ir(001) and Co/Ni/Ir(001)) is due to the fact that the data for 2Co/W(110) are presented for $E_{i}=4$ eV, where ExA is larger than SOA (see Fig. 7). Moreover, as seen in Fig. 8 at these wavevectors SOA is rather small. For the case of the Co/Ni/Ir(001) system the magnon nonreciprocity is presented for the incident energy of $E_{i}=10$ eV, where SOA is larger than ExA. This is also the case when looking at the $Q$-dependence of SOA, presented in Fig. 6.
@@
The results clearly demonstrate that both the incident beam energy and the choice of the substrate provide ways to tune the magnon nonreciprocity. This may inspire novel ideas for excitation of nonreciprocal ultrafast THz magnons for the application in the field of magnonics. We anticipate that the effect discussed here is not restricted to electron scattering experiments. It should also be observed in transport-based excitation schemes [61, 62, 63, 64].
Moreover, it is not limited to ultrathin films of $3d$ ferromagnets grown on metallic substrates. A similar effect is also expected when the underlying substrate is not metallic , e.g., dielectric oxides [65]. As long as the spin-orbit scattering is large one should be able to observe a magnon nonreciprocity. Analogously, the same effect is also expected for two-dimensional (2D) van der Waals magnets as well as heterostructures based on such 2D materials. One idea is to fabricate a bilayer made of a 2D ferromagnet and another 2D material which possesses a large SOC. Likewise, one may think of a 2D ferromagnet put on top of a topological insulator, e.g., Bi2Se3(0001) [14]. Since the interfacial electronic hybridization is not decisive for this phenomenon, one should, in principle, observe a notable magnon nonreciprocity.
In addition to ultrathin film systems discussed here, we anticipate that the phenomenon should also be present in other combinations of $3d$ magnetic elements and heavy-element substrates e.g., Pt. Other candidates are, the $3d$$5d$ alloys, e.g., FePt and CoPt or even more complex alloys, e.g., magnetic iridates and ruthenates, which possess a large SOC.
In principle, the effect shall also be present for systems with a moderate SOC. Since it scales with the strength of SOC, it can be best observed in the system in which this interaction is strong. Due to the complex nature of inelastic electron scattering it is not straightforward to estimate the expected magnon nonreciprocity for any arbitrary system. However, as a rule of thumb we expect that the effect would at least be proportional to the strength of SOC.
Another consideration is that our experiments were performed on high-quality epitaxial samples. In such a case the in-plane magnon momentum is well-defined and one can follow the dispersion relation of magnons with a high degree of accuracy. In the case of polycrystalline samples the magnon momentum is no longer a well-defined quantity and one deals with the magnons integrated over a certain range of momentum, depending on the degree of epitaxy. In order to observe the effect discussed here one requires a certain degree of epitaxy to avoid cancellation effects.
Since the magnon nonreciprocity is mainly caused by the fact that the excitation cross-section of magnons is different for the two opposite propagation directions (or the two opposite directions of $\bm{M}$), it is directly linked to the exchange process as the microscopic mechanism behind the magnon excitation. It is, therefore, expected that this kind of magnon nonreciprocity should be observed in all excitation schemes, which involve the exchange mechanism as the underlying mechanism for magnon excitation. For instance, in the transport-based tunneling experiments, in which a spin-polarized current is used to excite magnons e.g., in tunneling magnetoresistance type of devices (see for example [61, 62, 63, 64, 66, 67, 68]). Likewise, in the other excitation schemes, in which spin-polarized electrons are injected into a ferromagnetic layer, the observed magnon nonreciprocity should be present [69]. The effect should also be observed in other scattering methods based on inelastic neutrons and photons, as the amplitude of magnon wavepackets for opposite propagation directions (or the two opposite directions of $\bm{M}$) is intrinsically different.
If the aim is to utilize the effect in practical applications one needs to think of possible obstacles and possible strategies to overcome such obstacles. The feasibility of using THz magnons for information processing has been discussed in detail in Refs. [26, 25]. It is worth mentioning that one of the main advantages of these excitations is that they are ultrafast and are confined on nanometer length scales. Hence, they would allow for fabrication of nanoscale devices [70].
@@ Acknowledgements.
The financial support of the Deutsche Forschungsgemeinschaft (DFG) through the DFG Grants ZA 902/5-1, ZA 902/7-1 and ZA 902/8-1 is acknowledged. We would like to thank the Physikalisches Institut for hosting the group and for providing the infrastructure.